OpenStudy (anonymous):

what is the equation of a vertical line passing through the point (2,-3)

OpenStudy (anonymous):

x y - + - = 1 2 -3 3x - 2y = 6 -2y = 6 - 3x 2y = 3x - 6 or y = 3x/2 - 3

OpenStudy (anonymous):

Ok if it is a vertical line it will be in the form x - n where n is a real number As the line goes through a coordinate where the x value is given as 2, then the eqyuation wil be..... x = 2

OpenStudy (anonymous):

In first line it should read x=n